OpenStudy (sasogeek):

log2=0.3010 log3=0.4771 simplify log20 I'm stuck at log2+log2+log5, if i split up the log5, i would get log(2+3) but that isn't the same as log2+log3 so i'm probably missing something or using the wrong factors to start with... what am i supposed to do?

OpenStudy (dumbcow):

i think you are right, there is no way of getting log(20) in terms of log(3)

OpenStudy (sasogeek):

thanks for the confirmation, i thought i was lost

OpenStudy (mimi_x3):

But i think you can with \[\log(4*5) = \log4 + \log5\]

OpenStudy (sasogeek):

it's the same thing i expanded, but apparently you're supposed to use the values they gave you...

OpenStudy (mimi_x3):

ohh, but 2 and 3 won't work; so i thought of another approach..

OpenStudy (anonymous):

probably, u can use log 60 = log 2 + log 3 + log 10. It has the same meaning with log 60 = 2*log 2 + log 3 + log 5. so u can getting log 5 there. Compare log 60 = log 2 + log 3 + log 10.......(1) log 60 = 2*log 2 + log 3 + log 5.....(2) so log 2 + log 3 + log 10 = 2*log 2 + log 3 + log 5 log 5 = log 2 + log 3 + log 10 - 2*log 2 - log 3 then u can getting log 5 from here. log 20 = 2*log 2 + log 5 then substitute log 5, that we have had found. Perhaps, the way that i had given to u, can helping u to solve this problem. Thank you

OpenStudy (callisto):

log20 = log(2x10) = log2+log10 = log2+1 = 0.3010 +1 = 1.3010 then you can finish the task with the given value the a truth :)

OpenStudy (callisto):

it doesn't state that you must use the value log3 !

OpenStudy (anonymous):

wow, amazing. how fool i am. i was thinking too far. hehe

OpenStudy (anonymous):

20 = 2 * 10 -> log( 2 * 10) = log2 + log10 = .3010 + 1 = 1.3010